OpenStudy (anonymous):

Find the series solution of the differential equation:

OpenStudy (anonymous):

\[\large (6+x^2)y''-xy'+12y=0\] \[\large y=\sum_{n=0}^{\infty}a_nx^n,~~~y'=\sum_{n=1}^{\infty}na_nx^{n-1},~~~y''=\sum_{n=2}^{\infty}n(n-1)a_nx^{n-2}\] After plugging these in, I got \[\sum_{n=0}^{\infty}6(n+2)(n+1)a_{n+2}x^n+\sum_{n=2}^{\infty}n(n-1)a_nx^n-\sum_{n=1}^{\infty}na_nx^n+\sum_{n=0}^{\infty}12a_nx^n\] I was able to simplify this into: \[36a_3x+11a_1x+12a_2+12a_0+\]\[\sum_{n=2}^{\infty}[6(n^2+3n+2)a_{n+2}+(n^2-n)a_n-na_n+12a_n]x^n\] Where do I go after this?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

oops, I left out a n+2 in the second term exponent after I plugged them in, I believe

OpenStudy (anonymous):

Oh wait, no, I'm being stupid because I plugged in the x^2. :P

ganeshie8 (ganeshie8):

after rearranging you get just below right ? \[\sum_{n=0}^{\infty}[6(n^2+3n+2)a_{n+2}+(n^2-n)a_n-na_n+12a_n]x^n=0\]

OpenStudy (anonymous):

Plus the terms above it

ganeshie8 (ganeshie8):

how did you get them ?

OpenStudy (anonymous):

I plugged in n=0,1 to allow the indices to match without changing exponent degree

ganeshie8 (ganeshie8):

notice, below is true \[\sum\limits_{n=2}^5 n(n-1) =\sum\limits_{n=\color{red}{0}}^5 n(n-1) \]

OpenStudy (anonymous):

Oh, I see. So I guess I could then match the indices to n=1 instead of n=2

ganeshie8 (ganeshie8):

so you don't really need to pull anything out, simply change the index and you will be fine..

ganeshie8 (ganeshie8):

\[\sum_{n=0}^{\infty}6(n+2)(n+1)a_{n+2}x^n+\sum_{n=2}^{\infty}n(n-1)a_nx^n-\sum_{n=1}^{\infty}na_nx^n+\sum_{n=0}^{\infty}12a_nx^n\] is same as \[\sum_{n=0}^{\infty}6(n+2)(n+1)a_{n+2}x^n+\sum_{n=\color{red}{0}}^{\infty}n(n-1)a_nx^n-\sum_{n=\color{red}{0}}^{\infty}na_nx^n+\sum_{n=0}^{\infty}12a_nx^n\]

OpenStudy (anonymous):

I'm supposed to pull terms out. That's what I was taught in both the book and lecture. We have to do something with those terms to find a recursive formula that will help find the series solution.

ganeshie8 (ganeshie8):

Not in this problem, there is absolutely no need to pull anything out

OpenStudy (anonymous):

Okay, then how do I proceed with finding the solution?

ganeshie8 (ganeshie8):

the only reason for pulling few terms out is to align the exponents, in our case, all the exponents are aligning perfectly... so we're good.

OpenStudy (anonymous):

Wait why is the 3rd term in your post the same for both scenarios? Wouldn't the first term be a1x in the first scenario and 0 in the second?

OpenStudy (anonymous):

Oh but then 0+next term would be the same

ganeshie8 (ganeshie8):

Yes, the first term of \(\sum\limits_{n=0}^{\infty} na_n\) is 0 That's precicely the reason we don't need to pull anything out

OpenStudy (anonymous):

aha, gotcha. So how do I find the recursive formula for this guy?

ganeshie8 (ganeshie8):

\(\sum\limits_{n=1}^{\infty} na_nx^n =\sum\limits_{n=\color{red}{0}}^{\infty} na_nx^n \)

ganeshie8 (ganeshie8):

recursive relation is \[6(n^2+3n+2)a_{n+2} + (n^2-2n+12)a_n=0\]

ganeshie8 (ganeshie8):

it looks tough...

OpenStudy (anonymous):

Okay, so then I would solve for \(a_{n+2}\) right?

ganeshie8 (ganeshie8):

what exactly the instructions say ? do you really need to find the closed form or just finding the first few terms of power series is sufficient ?

ganeshie8 (ganeshie8):

because, solving above recurrence relation is really a pain...

OpenStudy (anonymous):

Our professor sent us home with the problem saying solve the differential equation with series solutions. In the past homework, we had to solve for the first 4-8 terms. So I imagine that that's what we have to do..

ganeshie8 (ganeshie8):

then it is simple, just isolate \(a_{n+2}\) and plugin n=0,1,2,...

OpenStudy (anonymous):

Okay, but then what's a0 and a1? Once I plug in n=0, I'll immediately start with a2

ganeshie8 (ganeshie8):

you may think of \(a_0\) and \(a_1\) as arbitrary constants and express all other coefficents in terms of \(a_0\) and \(a_1\)

OpenStudy (freckles):

Hey @CShrix I also got \[12(a_2+a_0)+(36a_3+11a_1)x \\ +\sum_{n=2}^{\infty} (6(n+2)(n+1)a_{n+2}+n(n-1)a_n-n a_n+12a_n)x^n=0 \\ \text{ I think this means we can say } \\ a_2=a_0 \text{ and } 36a_3=-11a_1 \text{ then plug in } n=2,3,4,5,... \text{ for the other part } \\ 6(n+2)(n+1)a_{n+2}+n(n-1)a_n-n a_n+12a_n=0\]

OpenStudy (freckles):

oops a_2=-a_0

ganeshie8 (ganeshie8):

@freckles \[12(a_2+a_0)+(36a_3+11a_1)x \\ +\sum_{n=2}^{\infty} (6(n+2)(n+1)a_{n+2}+n(n-1)a_n-n a_n+12a_n)x^n=0 \] is same as \[\sum_{n=\color{red}{0}}^{\infty} (6(n+2)(n+1)a_{n+2}+n(n-1)a_n-n a_n+12a_n)x^n=0 \] right?

OpenStudy (freckles):

no

OpenStudy (freckles):

but your way also works @ganeshie8 Like I plug in 0 and 1 and also got the a_2=-a_0 and 36a_3=-11a_1 into your thingy

OpenStudy (freckles):

oh is it the same as is what you said

OpenStudy (freckles):

I think so

OpenStudy (freckles):

But I was just trying to let @CShrix know what he had was also right

OpenStudy (anonymous):

Thanks @freckles !! I was scared that I was losing my mind because I thought I was wrong haha. Good to also see that there are other ways to think through the problem. Give me a few minutes so I can check and see if my terms are correct

OpenStudy (anonymous):

Okay, so for the first few terms, I have \(\large a_2=-a_0\) and \(\large a_3=\frac{-11}{36}a_1\) After plugging in values into the equation within the summation, I got: \[\large n=2~~~~~a_4=\frac{ 1 }{ 6 }a_0\]\[\large n=3~~~~~a_5=\frac{ 55 }{ 312 }a_1\]\[\large n=4~~~~~a_6=\frac{ -1 }{ 72 }a_0\]\[\large n=5~~~~~a_7=\frac{ -165 }{ 8736 }a_1\]

ganeshie8 (ganeshie8):

I did not check the arithmetic but it looks good to me!

OpenStudy (freckles):

maybe I did something wrong but I got \[a_5=\frac{-a_3}{8}=\frac{11}{36(8)}a_1\]

OpenStudy (freckles):

also I'm very prone to arithmetic mistakes right (I'm a little sick) or other mistakes

OpenStudy (freckles):

so you can double check that again

OpenStudy (anonymous):

Oh I think I made an error for that one. You're correct! Must've done something stupid lol

OpenStudy (anonymous):

So now that I have my terms, how can I plug it into a single combined solution?

OpenStudy (freckles):

you remember your solution form you chose a long time ago in the problem

OpenStudy (freckles):

\[y=\sum_{n=0}^{\infty} a_n x^n \\ =a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+ \cdots \]

OpenStudy (anonymous):

Ah, haha. Duh, that makes sense. I did that on the homework but I guess I had a memory lapse X) Thanks again to both @freckles and @ganeshie8 ! :)

OpenStudy (freckles):

i did check one more of your terms I got something different for a_6 too

OpenStudy (anonymous):

Yeah, I realized the algebraic mistake on that one too after I checked a_5 :P

OpenStudy (freckles):

k